### v2

# Assessing the replicability of RCTs in RWE emulations

Abstract
Abstract
Background: The standard regulatory approach to assess replication success is the two-trials rule, requiring both the original and the replication study to be significant with effect estimates in the same direction. The sceptical $p$-value was recently presented as an alternative method for the statistical assessment of the replicability of study results.
Methods: We review the statistical properties of the sceptical $p$-value and compare those to the two-trials rule. We extend the methodology to non-inferiority trials and describe how to invert the sceptical $p$-value to obtain confidence intervals. We illustrate the performance of the different methods using real-world evidence emulations of randomized, controlled trials (RCTs) conducted within the RCT DUPLICATE initiative.
Results: The sceptical $p$-value depends not only on the two $p$-values, but also on sample size and effect size of the two studies. It can be calibrated to have the same Type-I error rate as the two-trials rule, but has larger power to detect an existing effect. In the application to the results from the RCT DUPLICATE initiative, the sceptical $p$-value leads to qualitatively similar results than the two-trials rule, but tends to show more evidence for treatment effects compared to the two-trials rule.
Conclusion: The sceptical $p$-value represents a valid statistical measure to assess the replicability of study results and is especially useful in the context of real-world evidence emulations.
Keywords: Randomized clinical trial, Real-world evidence, Replication, Sceptical $p$-value, Two-trials rule.
1 Background
Randomized controlled trials (RCTs) represent the gold standard for proofing the efficacy of a new therapy (Byar et al., 1976; Abel and Koch, 1999; Concato et al., 2000). To reduce bias and guarantee a high internal validity of the results, an RCT is standardized to the highest degree and has strict in- and exclusion criteria. However, older, multi-morbid patients, women (particularly those lactating, pregnant, or able to become pregnant) or non-consenting patients are often excluded from RCTs (Van Spall et al., 2007; Food and Drug Administration, 2018) and as a result, a broad generalizability is sometimes missing (Schneeweiss and Avorn, 2005; Eichler et al., 2011). Thus, some vulnerable patient groups are under-represented in RCTs, resulting in a lack of evidence for treatment decisions.
Real-world evidence (RWE) plays an increasing role in research and for regulatory decision making, particularly due to an increasing number of potential data sources (US Food and Drug Administration, 2023; Franklin et al., 2019; Orsini et al., 2020). RWE is given as the clinical evidence based on so-called real-world data (RWD), i.e. data outside of RCTs given as data relating to patient health status or data routinely collected from different sources (Franklin et al., 2021). With the increase in purely virtual data sources (such as "digital twins" or "Metaverse©"), the term "RWD" is more and more criticized, as all study data come from the "real" world, leading to an increased use of the term "routine practice data" (Institute for Quality and Efficiency in Health Care, 2020). For the sake of clarity, however, the term RWD will be used as a synonym for routine practice data in the following and includes all type of data outside of RCTs.
Studies based on RWD can be used for post-market analysis of treatments in actual real-world care to ensure patient safety (Schneeweiss and Avorn, 2005; Eichler et al., 2011; Franklin et al., 2019), especially for research questions or patients groups that are difficult to address otherwise (Liu and Panagiotakos, 2022; Franklin and Schneeweiss, 2017; Franklin et al., 2019; Booth et al., 2019). In the literature known as efficacy-effectiveness gap, some studies reported a low performance of drugs and treatment in daily clinical practice (Bhandari et al., 2004; Eichler et al., 2011), but systematic analyses failed to proof a general significant difference between the effect estimate in observational studies and in RCTs (Concato et al., 2000; Golder et al., 2011; Kuss et al., 2011; Lonjon et al., 2014; Anglemyer et al., 2014). It is therefore of interest to determine if and when RCTs and well-designed RWE studies, closely aligned with the original RCT, reach the same conclusions. The RCT DUPLICATE initiative (Randomized, Controlled Trials Duplicated Using Prospective Longitudinal Insurance Claims: Applying Techniques of Epidemiology) consists of the emulation of 32 RCTs with RWD to check if and under which circumstances RWD studies explicitly designed to emulate RCTs are useful to draw the same causal conclusions (Franklin et al., 2021; Wang et al., 2023). A standardized and transparent study protocol for each RWE study was developed, with pre-defined patient selection and definition of primary analyses and study measures. Moreover, all emulation studies were a-priori registered before analyses were done (Franklin et al., 2021; Wang et al., 2023). All emulations were based on data of up to three different US claims data and propensity score matching with 120 pre-defined possible confounders was performed (Franklin et al., 2021; Wang et al., 2023).
It is common to distinguish direct and conceptual replications (Nosek and Errington, 2017). A direct replication is an attempt to reproduce a previously observed result based on new data, collected with a protocol as close as possible to the original study. Only the study sample size can be different and is often larger in replication studies to ensure sufficient power to confirm the original finding. In a conceptual replication, the aim is to show the robustness of a finding with a different study sample or study conditions. An emulation of an RCT is a conceptual replication, as it tries to confirm the result from an RCT with observational RWD (Wang et al., 2017). The possible problem of confounding through the lack of randomisation is usually addressed with statistical adjustments such as propensity score matching (Wang et al., 2023).
Whether direct or conceptual, the evaluation of the replicability of two studies (e.g. RCT vs replication given by the RWE emulation) is not straightforward and different statistical methods are available (Held, 2020; Errington et al., 2021). The regulatory agreement (also known as the two-trials rule, see e.g. Held, 2024) is a commonly used method to assess replicability of the two studies (Errington et al., 2021; Franklin et al., 2021). Regulatory agreement is fulfilled if the original effect and replication effect go in the same direction and are both significant. However, the pure assessment of the significance of the original trial and the replication study (two-trials rule – TTR) has the disadvantage that it does not directly take into account the effect sizes of the two studies (Held et al., 2022).
An alternative approach is to perform a meta-analysis of the RCT and RWE results. The two effect sizes are combined into an overall effect size and then significance is assessed. However, it is important to note that studies entering a meta-analysis are assumed to be interchangeable, which is sometimes a questionable assumption, since the studies were often conducted under different standards. In the case of the replication of an RCT with RWE in particular, this assumption is not fulfilled.
Another approach –  the so-called “sceptical” $p$-value –  was recently developed (Held, 2020; Held et al., 2022; Micheloud et al., 2023) and represents an alternative to the two-trials rule. It depends on the $p$-values of both studies and also on the variance ratio of both effect sizes. The original trial and the replication study are thus no longer interchangeable, a property which is particularly important with regards to the emulation of RCTs using RWD.
The aim of this paper is to investigate how the sceptical $p$-value performs in this context. To do so, the sceptical $p$-value will be used to evaluate the agreement of RWE emulation of RCTs given from the RCT DUPLICATE initiative (Franklin et al., 2021; Wang et al., 2023) and the results will be compared to the two-trials rule. First, we describe the data from the RCT DUPLICATE initiative in Section 2. Section 3 reviews the sceptical $p$-value and extends it to non-inferiority studies. Type-I error control and power for replication studies are discussed and confidence intervals for the combined effect are derived. In Section 4, the sceptical $p$-value is used to evaluate the replicability of emulations of RCTs given by the data set of the RCT DUPLICATE initiative (Franklin et al., 2021; Wang et al., 2023) and compared to the two-trials rule. The results are discussed in Section 5; especially in comparison with the conclusions from the original RCT DUPLICATE initiative.
2 Data source: RCT DUPLICATE initiative
For the study at hand, data of 32 RCTs and related RWE emulations from the RCT DUPLICATE initiative were available (Franklin et al., 2021; Wang et al., 2023). Detailed information about the RCT DUPLICATE initiative can be found in the literature (Franklin et al., 2020, 2021; Wang et al., 2023; Heyard et al., 2024). In short, published and ongoing RCTs were selected if the trials were relevant to regulatory decision making and were potentially replicable using RWD (Franklin et al., 2020, 2021). RCTs were considered to be potentially replicable if they fulfilled critical aspects concerning study protocol, primary outcome, inclusion and exclusion criteria. RCTs were chosen to get a mix of superiority and non-inferiority trials with large and small magnitudes of effect sizes and trials with active comparators as well as trials with active treatment and placebo added to active standard of care therapies (Franklin et al., 2020). All RCTs were required to be sufficiently large and well powered to guarantee reliable results.
For all emulation studies, the data of up to three different US claims data-sources were used to emulate the related RCTs: Optum Clinformatics (from 01/2004 to 03/2019), IBM Market Scan (from 2003 to 2017) and a subset of Medicare Part A, B and D (2011-2017 including all patients with a diabetes or heart failure diagnosis, 2009-2017 including all patients with a dispensation for an oral anticoagulant). Information about other insurance holders are not available for this database. As a result, 13 of the 32 RCTs were not emulated on the Medicare dataset (TRITON-TIMI, PLATO, ISAR-REACT5, TRANSCEND, ON-TARGET, HORIZON-PIVOTAL, VERO, P04334, D5896, IMPACT, POET-COPD and INSPIRE). The development of all emulation studies was standardized and transparent, with prospectively defined patient selection, primary analyses and registration (Franklin et al., 2021; Wang et al., 2023). Within a pre-defined time period, all available patient information in the claims data were included to the studies, with inclusion and exclusion criteria as closely as possible adopted from original RCTs. Moreover, a 1:1 nearest-neighbor propensity score matching with caliper width of 0.01 to control for more than 120 possible confounders was performed for all replication studies. Confounders were selected a priori and measured during 6 months before drug initiation. Detailed information about confounder selection can be found in Franklin et al. (2021).
For the analyses presented here, the pooled estimates from fixed effects meta-analysis of up to three different data sources were used for all emulation studies. These pooled estimates have also been analyzed in Heyard et al. (2024) with focus on differences between RCT and RWD effect size estimates due to design and population differences. For the original trials with a non-inferiority design, the margins were retrieved from the original protocols. For the sake of clarity, only original trials with hazard ratios (HRs) as the primary effect size were included, resulting in the exclusion of the LEAD2 trial due to a continuous outcome. However, in general, the sceptical $p$-value can also be used for continuous or binary outcomes. In addition, two further trials (ISAR-REACT5 and VERO trial) were excluded, because a meta-analysis was not performed due to significant heterogeneity between the estimates of the three US claims data detected by Cochran’s $Q$-test. Thus, $N=29$ pairs of original RCTs and related emulation studies (pooled data) were included for further analyses, $11$ with a superiority and $18$ with a non-inferiority design. The effect estimates of 28 trials were in the intended direction (HR < 1), out of which 26 were statistically significant. The two non-significant trials (INSPIRE and TRANSCEND) had one-sided $p$-values of $p=0.34$ and $p=0.10$, respectively. One trial (PRONOUNCE) had an effect estimate in the opposite direction (HR > 1), and is one of the two trials (with CAROLINA) for which the emulation was performed before the results of the RCT were made public. Hazard ratios with 95% confidence interval of all 29 analyzed trials from the RCT DUPLICATE initiative are presented in Fig. 1 and appendix table B.1.
Figure 1: Forest plot comparing RCT and RWE effect estimates (HR with 95% CI) from the RCT DUPLICATE initiative (Franklin et al., 2021; Wang et al., 2023). Individual non-inferiority margins were added for non-inferiority trials. Three trials were excluded due to a continuous outcome (LEAD2) and significant heterogeneity between the estimates of the three US claims data (ISAR-REACT5 and VERO).
3 Methods
3.1 The sceptical $p$-value
The lack of a standard approach to define replication success motivated Held (2020) to develop a new method: the sceptical $p$-value, which combines two principles: ‘the Analysis of Credibility’ by Matthews (2001a, b) and a prior-data conflict assessment proposed by Box (1980).
Assume that the effect estimate from the RCT is in the anticipated direction ($\mbox{HR}<1$, so $\hat{\theta}_{\mbox{\scriptsize RCT}}=\log(\mbox{HR})<0$) and is significant at level $\alpha$, i.e. one-sided $p$-value $p_{\mbox{\scriptsize RCT}}<\alpha$. This result from the RCT is challenged by a sceptical prior that renders the significant finding no longer convincing (lower limit of the posterior credible interval fixed at zero). The corresponding variance $\tau^{2}$ of the sceptical prior decreases with increasing (one-sided) $p$-value $p_{\mbox{\scriptsize RCT}}$, so RCTs with a larger $p$-value will lead to a wider sceptical prior. Note that the sceptical $p$-value for the alternative $\mbox{HR}<1$ is not available, if the effect estimate from the RCT is already in the opposite direction ($\mbox{HR}>1$), such as in the PRONOUNCE study.
In the second step, the sceptical prior needs to be compared with the RWE effect estimate $\hat{\theta}_{\mbox{\scriptsize RWE}}$ with standard error $\sigma_{\mbox{\scriptsize RWE}}$. Held (2020) proposed to assess prior-data conflict using the prior-predictive distribution of $\hat{\theta}_{\mbox{\scriptsize RWE}}$ (Spiegelhalter et al., 2004, Section 5.8). This results in the test statistic $t_{\mbox{\scriptsize Box}}={\hat{\theta}_{\mbox{\scriptsize RWE}}}/{\sqrt{\tau^{2}+\sigma_{\mbox{\scriptsize RWE}}^{2}}}$, which is compared against a standard normal $\operatorname{N}(0,1)$ distribution to obtain the tail probability $p_{\mbox{\scriptsize Box}}=\Phi(t_{\mbox{\scriptsize Box}})$ where $\Phi(.)$ denotes the standard normal cdf and smaller values of $p_{\mbox{\scriptsize Box}}$ indicate a larger conflict between the replication effect $\hat{\theta}_{\mbox{\scriptsize RWE}}$ and the sceptical prior. Replication success at level $\alpha$ is achieved if there is significant conflict between the prior and the RWE estimate, i.e. $p_{\mbox{\scriptsize Box}}\leq\alpha$.
The approach needs some adaptions for non-inferiority trials. If $\delta$ denotes the non-inferiority margin on the log hazard ratio scale, then the sceptical prior is now centered at $\delta$. We then obtain the modified test statistic
$t_{\mbox{\scriptsize Box}}=({\hat{\theta}_{\mbox{\scriptsize RWE}}-\delta})/{\sqrt{\tau^{2}+\sigma_{\mbox{\scriptsize RWE}}^{2}}}.$ (1)
3.1.1 The original formulation
The value of $p_{\mbox{\scriptsize Box}}$ depends on the significance level $\alpha$, thus rendering the interpretation difficult. The sceptical $p$-value $p_{S}$ is the smallest level $\alpha$ for which replication success can be achieved, so no longer depends on $\alpha$. Some algebra shows that the sceptical $p$-value depends on the (one-sided) original and replication $p$-value $p_{\mbox{\scriptsize RCT}}$ and $p_{\mbox{\scriptsize RWE}}$ and the variance ratio $c=\sigma_{\mbox{\scriptsize RCT}}^{2}/\sigma_{\mbox{\scriptsize RWE}}^{2}$, here $\sigma_{\mbox{\scriptsize RCT}}$ denotes the standard error of the RCT effect estimate. For non-inferiority trials, the $p$-values are calculated for the null hypothesis that the true effect is equal to the non-inferiority margin $\delta$.
The sceptical $p$-value can be interpreted quantitatively: Smaller $p_{S}$ values indicate a higher degree of replication success. The sceptical $p$-value is one-sided to ensure that replication can only occur if the replication effect estimate has the same sign as the original effect estimate. Further properties of the sceptical $p$-value are studied in Held (2020). First, the approach can be shown to be more stringent than the two-trials rule: if replication success is achieved ($p_{S}\leq\alpha$), then two-trials rule is also fulfilled ($p_{\mbox{\scriptsize TTR}}=\max\{p_{\mbox{\scriptsize RCT}},p_{\mbox{\scriptsize RWE}}\}\leq\alpha$) but not necessarily the other way around. As a consequence, the overall Type-I error rate of the sceptical $p$-value (assuming that both the original and the replication effect estimate is zero) is smaller than the Type-I error rate $\alpha^{2}$ of the two-trials rule. Second, for fixed original and replication $p$-value, the sceptical $p$-value increases with increasing variance ratio $c$. The variance ratio can be rewritten as the relative sample size $c=n_{\mbox{\scriptsize RWE}}/n_{\mbox{\scriptsize RCT}}$ (replication to original), if we assume that the unit variance in both studies is the same. This implies that a replication study with a large sample size will show a smaller degree of replication success than a replication study with the same replication $p$-value, but a smaller sample size. The method hence requires replication studies with large sample sizes to be more convincing than those with small sample sizes to achieve the same degree of replication success.
3.1.2 Exact Type-I error control
The sceptical $p$-value has been recently re-calibrated to achieve exact overall Type-I error control across both studies (Micheloud et al., 2023) for every value of $c$. The new version has still improved project power compared to the two-trials rule but can now lead to replication success even if one of the two studies is not significant. In what follows we will report the re-calibrated "controlled" version of the sceptical $p$-value.
The exact overall calibration at level $\alpha^{2}$ implies that we can invert the sceptical $p$-value to obtain one-sided confidence intervals for the combined effect estimate. The upper limit $\theta_{u}$ of a 97.5% CI for the log hazard ratio is therefore obtained as the value of $\delta=\theta_{u}$, where the modified test statistic (1) gives a squared sceptical $p$-value of 0.025. Exponentiation gives an upper limit on the hazard ratio scale, which can directly be compared to the non-inferiority (or superiority) margin. This approach provides an alternative to standard meta-analytic pooling, which may not be compatible to the sceptical $p$-value nor the two-trials rule.
3.1.3 A comparison
The (controlled) sceptical $p$-value has exact overall Type-I error control (across both studies) at level $\alpha^{2}$ for all values of the variance ratio $c$. The two-trials rule also offers exact Type-I error control at level $\alpha^{2}$. However, the corresponding $p$-value $p_{\mbox{\scriptsize TTR}}=\max\{p_{\mbox{\scriptsize RCT}},p_{\mbox{\scriptsize RWE}}\}$ cannot be smaller than $p_{\mbox{\scriptsize RCT}}$ nor $p_{\mbox{\scriptsize RWE}}$. Figure 2 compares the behaviour of the sceptical $p$-value $p_{S}$ and the corresponding $p$-value $p_{\mbox{\scriptsize TTR}}$ of the two-trials rule as a function of the relative sample size for fixed original $p$-values. The calculations assume that the relative effect size $\hat{\theta}_{\mbox{\scriptsize RWE}}/\hat{\theta}_{\mbox{\scriptsize RCT}}$ is fixed at 1 (top) and 0.75 (bottom), respectively. Note that for fixed relative effect size, the replication $p$-value depends on the assumed relative effect size and the relative sample size.
Figure 2 shows that the sceptical $p$-value behaves very similar to the two-trials rule if the relative sample size is small. However, it decreases smoothly towards zero and can get smaller than $p_{\mbox{\scriptsize RCT}}$ if the relative sample size $c$ becomes large, whereas the two-trials rule $p$-value has an edge and then stays constant, as it cannot get smaller than $p_{\mbox{\scriptsize RCT}}$.
Figure 2: Sceptical $p$-value (square root) and $p$-value from the two-trials rule shown as a function of relative sample size (RWE to RCT) for fixed $p$-value of the original study (dashed grey line) and fixed relative effect size (first row: 1, second row: 0.75).
3.2 Power of the replication study
The power of the replication study, i.e. the probability that a replication study with a certain sample size leads to replication success, has been extensively studied (see for example Anderson and Maxwell, 2017; Anderson and Kelley, 2022; van Zwet and Goodman, 2022; Micheloud and Held, 2022; Micheloud et al., 2023). The power is usually calculated as the probability to reach a significant replication result conditional on the effect estimate from the original trial. This concept is illustrated in Figure 3 using the TRITON-TIMI trial from the RCT DUPLICATE initiative. The conditional power of the two-trials rule at the one-sided level $\alpha=0.025$ is plotted as a function of the effect estimate from the RCT (log hazard ratio) assuming that the variance is the same in the original trial and the emulation study ($c=1$). The original effect estimate in the TRITON-TIMI example is $\hat{\theta}_{\mbox{\scriptsize RCT}}=-0.21$, so with $c=1$, a conditional power of 98.1% is achieved. However, the original effect estimate comes with uncertainty, represented by the density in the top axis, and the 95% confidence interval $[-0.31,-0.11]$ around the original effect estimate. The conditional power for the values inside this confidence interval greatly vary, from 52.6% to 100%. One way to incorporate this uncertainty is to compute the predictive power (Spiegelhalter and Freedman, 1986) by averaging the conditional power with respect to the distribution of the original effect estimate. The density of the conditional power is shown in the right axis of Figure 3 and the average is 92.6%, so considerably smaller than 98.1%.
Figure 3: Conditional power as a function of the original effect estimate. The original effect estimate $\hat{\theta}_{\mbox{\scriptsize RCT}}=-0.21$ of the TRITON-TIMI trial and the corresponding conditional power of 98.1% are colored in red. The grey lines in the top and right axis are the distributions of the original effect estimate and the corresponding conditional power, respectively. The average conditional power of 92.6% (in blue) is the predictive power. The green intervals represent the range given by the 95% confidence interval of the original effect estimate.
If the power to detect the effect from the original study is not too large (< 95%), the conditional Type-I error rate (the probability of replication success if the replication effect estimate is zero) of the sceptical $p$-value is shown to be bounded and never larger than $2\alpha$ (Micheloud et al., 2023, Section 3.4). Such a two-fold increase in Type-I error rate is the price to be paid for increased power, but considered to be acceptable even in the regulatory setting (Rosenkranz, 2002).
Power calculations are also used to determine which sample size is required in the replication study to reach a certain power, usually between 80% and 90%. While replication sample size calculation is not relevant in the RCT DUPLICATE study, where all available patient information in the claims data was used, it is of interest to know the power of a RWE study with the available data. It is therefore advisable to determine the power for an existing dataset before calculating the effect estimates for the emulation, to ensure that the data are indeed sufficient. Conversely, calculating the minimum necessary sample size can also be beneficial in the context of RWE, for example, to understand the cohort size required after propensity score matching. If the power is insufficient or if the sample size is inadequate after matching, the chosen dataset is likely not suitable for the planned emulation. In Section 4, we calculate the conditional and predictive power of the two-trials rule and of the sceptical $p$-value for all the RWE emulations (Micheloud and Held, 2022; Micheloud et al., 2023).
4 Results
4.1 Replication of RCTs
Results of the replication success assessment with the two methods are presented in Table 1. The variance ratio $c$ is larger than 10 for two studies (TECOS, INSPIRE), see also Figure 1. TECOS is a non-inferiority study, where both the RCT and the RWE study showed non-inferiority with respect to a HR margin of 1.4. INSPIRE is a superiority study, where both effect estimates have been close to 1, but the RWE estimate is slightly smaller and more much precise. The sceptical $p$-value is smaller or equal to the $p$-value from the two-trials rule for the vast majority of the studies (27/29 studies). Note also that in three cases (CARMELINA, TRANSCEND, INSPIRE) the sceptical $p$-value is smaller than $p_{\mbox{\scriptsize RCT}}$ and in seven cases (TRITON-TIMI, PLATO, HORIZON-PIVOTAL, DAPA-CKD, P04334, D5896, and PRONOUNCE) it is smaller than $p_{\mbox{\scriptsize RWE}}$, both is impossible for $p_{\mbox{\scriptsize TTR}}$.
At the one-sided level $\alpha=0.025$, the same conclusion is drawn with the two-trials rule and the sceptical $p$-value: The same 20 out of 29 (69%) emulations successfully replicate the corresponding original RCT results. However, it was noticeable that more studies failed to replicate the original effect, if no Medicare data were available for the pooled effect. If Medicare data were available, 16 of 19 (84%) emulations successfully replicated the original effect, compared to the lower rate (5 of 10, 50%) of successfully replications, when Medicare data was not available. This finding was independent from the used metric, i.e. the sceptical $p$-value and the $p$-value from the two-trials rule lead to the same rates of success.
Study $p_{\mbox{\scriptsize RCT}}$ $p_{\mbox{\scriptsize RWE}}$ $c$ $p_{\mbox{\scriptsize TTR}}$ $p_{S}$
1 LEADER < 0.0001 < 0.0001 2.6 < 0.0001 < 0.0001
2 DECLARE < 0.0001 < 0.0001 0.7 < 0.0001 < 0.0001
3 EMPA-REG < 0.0001 < 0.0001 1.2 < 0.0001 < 0.0001
4 CANVAS < 0.0001 < 0.0001 1.8 < 0.0001 < 0.0001
5 CARMELINA 0.0003 < 0.0001 4.2 0.0003 < 0.0001
6 TECOS < 0.0001 < 0.0001 14.3 < 0.0001 < 0.0001
7 SAVOR-TIMI < 0.0001 < 0.0001 3.5 < 0.0001 < 0.0001
8 TRITON-TIMI < 0.0001 0.007 1.0 0.007 0.003
9 PLATO < 0.0001 0.056 0.7 0.056 0.031
10 ARISTOTLE < 0.0001 < 0.0001 2.7 < 0.0001 < 0.0001
11 RE-LY < 0.0001 < 0.0001 1.2 < 0.0001 < 0.0001
12 ROCKET-AF < 0.0001 < 0.0001 2.2 < 0.0001 < 0.0001
13 EINSTEIN-DVT < 0.0001 < 0.0001 5.3 < 0.0001 < 0.0001
14 EINSTEIN-PE 0.0018 < 0.0001 4.4 0.0018 < 0.0001
15 RECOVER II 0.0002 < 0.0001 1.4 0.0002 0.0002
16 AMPLIFY < 0.0001 < 0.0001 0.7 < 0.0001 < 0.0001
17 RECORD1 < 0.0001 < 0.0001 1.3 < 0.0001 < 0.0001
18 TRANSCEND 0.10 0.002 2.3 0.10 0.064
19 ON-TARGET 0.001 < 0.0001 0.9 0.001 < 0.0001
20 HORIZON-PIVOTAL 0.001 0.014 1.8 0.014 0.01
21 DAPA-CKD < 0.0001 0.16 0.2 0.16 0.15
22 PARADIGM-HF < 0.0001 0.63 0.6 0.63 0.65
23 P04334 < 0.0001 0.015 1.2 0.015 0.004
24 D5896 0.002 0.046 1.0 0.046 0.03
25 IMPACT < 0.0001 1.00 0.5 1.00 1.00
26 POET-COPD < 0.0001 0.66 0.7 0.66 0.68
27 INSPIRE 0.34 < 0.0001 19.9 0.34 0.25
28 CAROLINA 0.0001 < 0.0001 1.2 0.0001 < 0.0001
29 PRONOUNCE 0.73 0.95 4.7 0.95 NA
Table 1: One-sided $p$-values $p_{RCT}$ of the original trial and $p_{RWE}$ of the emulation, as well as variance ratio $c$ and one-sided $p$-values $p_{\mbox{\scriptsize TTR}}$ of the two-trials rule and $p_{S}$ of the sceptical $p$-value.
Figure 4 shows $\hat{\theta}_{\mbox{\small RCT}}$ log(margin) versus $\hat{\theta}_{\mbox{\small RWE}}$ log(margin) for the 29 study pairs in the RCT DUPLICATE initiative. The margin is $1$ in the superiority studies. Points below the diagonal represent cases where the effect estimate from the RWE study is shrunken as compared to the effect estimate from the original trial. This shrinkage is towards 00 for superiority studies and towards the margin for non-inferiority studies. Replication effect estimates usually present a large amount of shrinkage as compared to their original counterpart in recent large-scale replication projects (e.g. in Open Science Collaboration, 2015; Camerer et al., 2016, 2018). This is not the case here.
Figure 4: $\hat{\theta}_{\mbox{\small RCT}}$ log(margin) versus $\hat{\theta}_{\mbox{\small RWE}}$ log(margin) with their respective 95% confidence interval for the studies in the RCT DUPLICATE initiative (without RECORD1, where differences are larger).
4.2 Replication power
Results are displayed in Table 2. The average conditional power (Amrhein et al., 2019; McShane et al., 2020) is 85.9% with the two-trials rule and 87.0% with the sceptical $p$-value. These values reduce to 83.5%, and 85.0%, respectively, for predictive power. If the true underlying effect is the same in the original trial and in the emulation study and both were conducted with high standards, we would expect the average predictive power to be equal to the proportion of success, which is smaller in the RCT DUPLICATE study (69% with both methods). The proportion of replication success as well as the average predictive power in the RCT DUPLICATE initiative for other values of $\alpha$ are shown in Figure A.1.
The three RCTs with a non-significant original $p$-value (PRONOUNCE, INSPIRE and TRANSCEND) have a power of 00 with all three methods. By definition, the power based on a non-significant original trial is always 00 with the two-trials rule. There is no such requirement on $p_{\mbox{\scriptsize RCT}}$ for the sceptical $p$-value but large relative sample sizes are required to reach an acceptable power level for non-significant original studies.
For the 26 other studies, the power of the two-trials rule and the sceptical $p$-values is very similar, with slightly higher values with the sceptical $p$-values. This is because the original $p$-values are relatively small in every case, which is favored by the sceptical $p$-values (see Held et al., 2022; Micheloud et al., 2023, Sections 3.1 and  3.4, respectively).
4.3 Confidence intervals
We have also calculated confidence intervals for the hazard ratio based on inversion of the sceptical $p$-value and compared those to the ones obtained from applying a fixed effect meta-analysis to the results from the RCT and RWD. The results are given in Table B.1 in Appendix B. The upper limit of both confidence intervals are similar, the upper limit of the sceptical CI being often slightly larger than the meta-analytic one. The differences get larger when the uncertainty of the effect estimates is substantially different, for example for DAPA-CKD or INSPIRE, where the meta-analytic pooled effect estimate is dominated by the larger study. This is also the case for the three superiority studies PARADIGM-HF, IMPACT and POET-COPD, where the meta-analytic confidence interval indicates significance (at the standard two-sided 95% level) with an upper CI limit below 1, although the RWE effect estimate is above 1 in all three cases. In contrast, the upper limit of the sceptical CI is always above one, in line with the large sceptical $p$-values for those studies reported Table 1. This illustrates that meta-analytic pooling can still flag a positive (significant) finding, even if one of the two studies has an effect in the wrong direction (Held, 2024).
Conditional power Predictive power
Study $p_{\mbox{\scriptsize RCT}}$ two-trials rule sceptical $p$ two-trials rule sceptical $p$
1 LEADER < 0.0001 100.0 100.0 100.0 100.0
2 DECLARE < 0.0001 100.0 100.0 99.7 99.9
3 EMPA-REG < 0.0001 100.0 100.0 99.8 99.9
4 CANVAS < 0.0001 100.0 100.0 100.0 100.0
5 CARMELINA 0.0003 100.0 100.0 98.8 99.3
6 TECOS < 0.0001 100.0 100.0 100.0 100.0
7 SAVOR-TIMI < 0.0001 100.0 100.0 99.9 99.9
8 TRITON-TIMI < 0.0001 98.1 99.2 92.6 95.3
9 PLATO < 0.0001 91.2 95.0 84.8 89.3
10 ARISTOTLE < 0.0001 100.0 100.0 100.0 100.0
11 RE-LY < 0.0001 100.0 100.0 100.0 100.0
12 ROCKET-AF < 0.0001 100.0 100.0 100.0 100.0
13 EINSTEIN-DVT < 0.0001 100.0 100.0 100.0 100.0
14 EINSTEIN-PE 0.0018 100.0 100.0 100.0 97.0
15 RECOVER II 0.0002 98.7 99.4 92.4 94.9
16 AMPLIFY < 0.0001 95.2 97.6 90.1 93.7
17 RECORD1 < 0.0001 100.0 100.0 100.0 100.0
18 TRANSCEND 0.10 0.0 0.0 0.0 0.0
19 ON-TARGET 0.001 80.5 85.6 73.4 77.9
20 HORIZON-PIVOTAL 0.001 98.0 98.9 89.3 91.8
21 DAPA-CKD < 0.0001 59.1 65.6 58.5 64.6
22 PARADIGM-HF < 0.0001 97.3 98.7 93.5 96.1
23 P04334 < 0.0001 99.9 100.0 98.2 99.1
24 D5896 0.002 81.2 85.7 73.5 77.5
25 IMPACT < 0.0001 96.7 98.3 93.4 95.9
26 POET-COPD < 0.0001 97.6 98.9 93.4 96.0
27 INSPIRE 0.34 0.0 0.0 0.0 0.0
28 CAROLINA 0.0001 97.3 98.8 90.6 93.7
29 PRONOUNCE 0.73 0.0 0.0 0.0 0.0
30 Average 85.9 87.0 83.5 85.0
Table 2: Conditional and predictive power (in %) at the one-sided level $\alpha=0.025$ with the two-trials rule and the sceptical $p$-value.
5 Discussion
A target trial emulation is a conceptual emulation of a "hypothetical" RCT, using a two-step process (Hernán and Robins, 2016; Hernán et al., 2022). In the first step, the causal questions are defined in a protocol for the hypothetical RCT, which has to specify the key elements to define the causal estimands (i.e. eligibility criteria, treatment strategies, treatment assignment, the start and end of follow-up, outcomes, causal contrasts) and the data analysis plan (Hernán and Robins, 2016). The RCT described in the protocol is defined as the target study for causal inference, which will be emulated using observational data in the second step (Hernán and Robins, 2016; Hernán et al., 2022). Furthermore, several studies used an existing RCT instead of a hypothetical RCT, which will be emulated by the target trial emulation (Deng et al., 2022; Merola et al., 2022; Leening and Boersma, 2024; Walker et al., 2024; Petito et al., 2020; McGrath et al., 2021; Khosrow-Khavar et al., 2022; Gallivan et al., 2023; Yoon et al., 2023). However, even with the existing RCT as a template for the target trial emulation, the effect estimates of the RWD emulation are not always compared with those of the original trial (Deng et al., 2022; Petito et al., 2020; McGrath et al., 2021; Khosrow-Khavar et al., 2022; Gallivan et al., 2023). A statistical comparison of the emulation with the original RCT further substantiates the results and increase the resulting evidence. As seen in the data presented, the sceptical $p$-value can be an excellent extension to evaluate the agreement and thus complements the evidence from target trial emulations of existing RCTs.
The emulations of the 32 evaluated RCTs were previous implemented and analyzed by the RCT DUPLICATE Initiative (Wang et al., 2023). In this project, Wang et al. (2023) used different binary agreement metrics to check the replication success of the 32 emulations using claims data of three different US insurance companies. First, named full statistical significance agreement, is fulfilled, if the estimates (log(HR)) and the related 95% CIs are on the same side of the null (without consideration of the predefined non-inferiority margins in the case of non-inferiority trials). 23 (72%) of the 32 emulations met statistical significance agreement (Wang et al., 2023). This binary agreement metric should be equivalent to the two-trials rule. However, for two studies, i.e. D5896 and PRONOUNCE, the $p$-value of the two-trials rule delivered a different conclusion of the replication success. Wang et al. (2023) reported that statistical agreement was fulfilled for all of the two emulations, whereby in our analysis, $p_{\mbox{\scriptsize TTR}}>0.025$ holds. In the case of PRONOUNCE, the original RCT had a non-significant treatment effect, resulting in an non-significant $p_{\mbox{\scriptsize TTR}}$ due to the definition of the $p$-value. Moreover, the formalism of $p_{\mbox{\scriptsize TTR}}$ is not meaningful in the case of non-significant effects of the original trial. In the case of the D5896 trial, Wang et al. (2023) did not use the non-inferiority margins for the definition of full statistical agreement, as it was simply defined as having point estimates and confidence intervals on the same side of null. Using sceptical $p$-value, it is possible to account for the non-inferiority margins, which is one major advantage of this formalism compared to the binary agreement metrics.
As the second binary agreement metric, Wang et al. (2023) analysed estimated agreement, whereby the estimate of the emulation has to be an element of the 95% CI of the original effect, which was met for 66% of the emulations. As the last one, Wang et al. (2023) evaluated, if the absolute standardized differences of the estimates are smaller then 1.96, which was fulfilled in 72% of the emulations. Using the sceptical $p$-value, 69% of the emulations were evaluated as successful replications. Beside the emulations with non-significant original effects, the findings of the study at hand are in accordance to findings of Wang et al. (2023).
However, in general, it is not always possible to replicate the results of an RCT, especially if there are fundamental design differences between the RCT and RWE (Wang et al., 2023; Heyard et al., 2024). Wang et al. (2023) stated that the conceptual emulation of the original RCT leads to an independent study that emulate the design of the original trial instead of replicating the population of the RCT. Potential differences of the two population (e.g. different distribution of comorbidities) may result in differences in the effect size estimates (Wang et al., 2023). Moreover, Heyard et al. (2024) analyzed differences between the original RCTs and the related RWE emulations using meta-regression models. They concluded that most of the heterogeneity between RCT and RWD emulation could be explained by delayed effect of treatment, discontinuation of treatment during run-in period, and treatment started in hospital (Heyard et al., 2024). The sceptical $p$-value itself cannot distinguish between replication failure due to design differences, incorrect model specification or differences due to population differences. Instead, the sceptical $p$-value evaluates and quantifies the agreement between two effect estimators. A successful replication of an RCT using RWD, indicated by a significant sceptical $p$-value, suggests the external validity of the RCT results. Conversely, if the replication is unsuccessful, the p-value does not provide insight into the reasons. These issues must be addressed using other methods. In studies based on RWD, it is essential to adequately address data heterogeneity methodologically, for example, using propensity score methods. An important aspect is achieving sufficient balance between the treatment groups. However, it is possible that the treatment effect observed in an RCT cannot be replicated in RWD due to fundamentally different cohorts. This is less a failure of RWE and may indicate that external validity is limited.
There are other studies available that emulate RCTs using RWD, where the binary metrics regulatory agreement, estimated agreement and standardized difference agreement were often used to evaluate the agreement of the effect estimates from the original trials and the related emulation study (Jang et al., 2022; Jin et al., 2022; Merola et al., 2023; Antoine et al., 2024; Signori et al., 2024; D’Andrea et al., 2024). However, the two-trials rule and the sceptical $p$-values are good measures to evaluate the agreement between an original RCT and the corresponding emulation study and have benefits compared to binary metrics. Besides the advantages discussed above, both methods allow to calculate the power of the emulation study based on the effect estimate found in the RCT, with and without uncertainty incorporation. For all 29 studies considered here, the conditional and predictive power is larger with the sceptical $p$-values than with the two-trials rule. Furthermore, the replication rate is in general slightly higher with the sceptical $p$-value.
In this paper, we focused on the controlled version of the sceptical $p$-value, to ensure a fair comparison with the two-trials rule in terms of overall Type-I error control. An alternative to the controlled re-calibration is the golden re-calibration of the sceptical $p$-value (Held et al., 2022). The golden version is less stringent than the original formulation, because, to establish replication success, original trial and replication study do not both necessarily need to be significant at level $\alpha$. However, a borderline significant original trial ($p_{\mbox{\scriptsize RCT}}=\alpha$) cannot lead to replication success, if there is shrinkage of the replication effect estimate. As shrinkage of the effect estimate from the emulation study as compared to effect estimate from the original trial seems to not be an issue in this context, one could favor the controlled sceptical $p$-value over the golden sceptical $p$-value. Moreover, a comparison of the golden version with the two-trials rule is more difficult, as the two methods have different Type-I error rates.
The presented work had several limitations. No formalism to assess the replication success of an original null effect using the sceptical $p$-value was available (Pawel et al., 2024). Recently Micheloud and Held (2024) provided an extension of the sceptical $p$-value to equivalence studies. However, a trial that was initially planned for superiority, but fails to show a significant treatment effect, should not be interpreted as an equivalence study and, thus, the formalism for equivalence studies should not be assigned in this case. Instead of performing a second study with the wrong design, a new study with the aim of proofing the equivalence should be initialized (and then replicated). Moreover, as stated above, fundamental design differences between RCT and RWD emulation may result in bias, which will also affect the effect estimate of the RWD study.
6 Conclusion
The sceptical $p$-value represents a valid statistical measure to evaluate the replicability of the study results, especially useful in the context of real-world evidence.
7 List of abbreviations
Abbreviation Explanation
CI Confidence interval
HR Hazard ratio
RCT Randomized controlled trial
RWD Real world data
RWE Real world evidence
TTR Two-trials rule
8 Declarations
8.1 Ethics approval and consent to participate
Not applicable.
8.2 Consent for publication
Not applicable.
8.3 Availability of data and materials
The R package ReplicationSuccess, available on CRAN at https://CRAN.R-project.org/package=ReplicationSuccess, has been used for the computation of the sceptical $p$-value and all power calculations. The data and the code to reproduce all the Figures and Tables can be found at https://github.com/CharlotteMicheloud/RWE.
8.4 Competing interests
JK reports research funding from the German Society for Trauma Surgery sponsored by Stryker, outside the submitted work. SE, CM, RH and LH have no conflict of interest to declare.
8.5 Funding
None.
8.6 Authors’ contributions
Jeanette Köppe: Conceptualization, Methodology, Formal analysis, Software, Writing - Original draft, Writing - Review & Editing, Visualization Charlotte Micheloud: Methodology, Formal analysis, Software, Writing - Original draft, Writing - Review & Editing, Visualization Stella Erdmann: Conceptualization, Methodology, Formal analysis, Writing - Review & Editing Rachel Heyard: Data curation, Writing - Review & Editing Leonhard Held: Conceptualization, Methodology, Formal Analysis, Writing - Original draft, Writing - Review & Editing, Visualization
8.7 Acknowledgements
We thank the RCT DUPLICATE team for their effort and for making their data publicly available.
Appendix A Replication power
Figure A.1 shows the proportion of replication success as well as the average predictive power as a function of the one-sided level $\alpha$. The proportion of success is for some values of $\alpha$ slightly smaller with the two-trials rule than with the sceptical $p$-value, but the two methods are in general very close to each other.
Figure A.1: Proportion of success and average predictive power as a function of the one-sided level $\alpha$.
Appendix B Confidence intervals for combined effects
Study NI RCT pooled RWE meta analysis sceptical $p$-value
margin HR (95% CI) HR (95% CI) HR (95% CI) (one-sided 97.5% CI)
1 LEADER 1.30 0.87 (0.78; 0.97) 0.82 (0.76; 0.87) 0.83 (0.79; 0.88) (0; 0.91)
2 DECLARE 1.30 0.83 (0.73; 0.95) 0.69 (0.59; 0.81) 0.77 (0.70; 0.85) (0; 0.88)
3 EMPA-REG 1.30 0.86 (0.74; 0.99) 0.83 (0.73; 0.95) 0.84 (0.76; 0.93) (0; 0.92)
4 CANVAS 1.30 0.86 (0.75; 0.97) 0.77 (0.70; 0.85) 0.80 (0.74; 0.87) (0; 0.91)
5 CARMELINA 1.30 1.02 (0.89; 1.17) 0.90 (0.84; 0.96) 0.92 (0.87; 0.98) (0; 1.07)
6 TECOS 1.30 0.98 (0.88; 1.09) 0.89 (0.86; 0.91) 0.90 (0.87; 0.92) (0; 1.01)
7 SAVOR-TIMI 1.30 1.00 (0.89; 1.12) 0.81 (0.76; 0.86) 0.85 (0.80; 0.90) (0; 1.04)
8 TRITON-TIMI 1.00 0.81 (0.73; 0.90) 0.88 (0.79; 0.97) 0.84 (0.79; 0.91) (0; 0.92)
9 PLATO 1.00 0.84 (0.77; 0.92) 0.92 (0.83; 1.02) 0.87 (0.82; 0.93) (0; 0.96)
10 ARISTOTLE 1.38 0.79 (0.66; 0.95) 0.68 (0.61; 0.76) 0.71 (0.64; 0.78) (0; 0.85)
11 RE-LY 1.46 0.66 (0.53; 0.82) 0.73 (0.60; 0.90) 0.70 (0.60; 0.81) (0; 0.80)
12 ROCKET-AF 1.46 0.79 (0.66; 0.96) 0.70 (0.62; 0.80) 0.73 (0.65; 0.81) (0; 0.85)
13 EINSTEIN-DVT 2.00 0.68 (0.44; 1.04) 0.75 (0.62; 0.90) 0.74 (0.62; 0.88) (0; 0.86)
14 EINSTEIN-PE 2.00 1.12 (0.75; 1.64) 0.67 (0.55; 0.80) 0.74 (0.62; 0.87) (0; 1.29)
15 RECOVER II 2.75 1.08 (0.64; 1.80) 1.15 (0.74; 1.78) 1.12 (0.80; 1.57) (0; 1.48)
16 AMPLIFY 1.80 0.84 (0.60; 1.18) 0.81 (0.54; 1.23) 0.83 (0.64; 1.07) (0; 1.03)
17 RECORD1 1.95 0.25 (0.14; 0.47) 0.17 (0.10; 0.29) 0.20 (0.13; 0.30) (0; 0.33)
18 TRANSCEND 1.00 0.92 (0.81; 1.05) 0.88 (0.81; 0.96) 0.89 (0.83; 0.96) (0; 0.97)
19 ON-TARGET 1.13 1.01 (0.94; 1.09) 0.83 (0.77; 0.90) 0.92 (0.87; 0.97) (0; 1.04)
20 HORIZON-PIVOTAL 1.00 0.59 (0.42; 0.83) 0.75 (0.58; 0.97) 0.69 (0.56; 0.84) (0; 0.84)
21 DAPA-CKD 1.00 0.61 (0.51; 0.72) 0.80 (0.52; 1.26) 0.63 (0.54; 0.74) (0; 0.99)
22 PARADIGM-HF 1.00 0.80 (0.73; 0.87) 1.02 (0.91; 1.14) 0.88 (0.82; 0.94) (0; 1.07)
23 P04334 1.00 0.56 (0.44; 0.72) 0.78 (0.62; 0.97) 0.67 (0.57; 0.79) (0; 0.86)
24 D5896 2.00 1.07 (0.70; 1.65) 1.38 (0.90; 2.13) 1.21 (0.90; 1.65) (0; 1.68)
25 IMPACT 1.00 0.85 (0.80; 0.90) 1.13 (1.04; 1.23) 0.93 (0.89; 0.98) (0; 1.17)
26 POET-COPD 1.00 0.83 (0.77; 0.90) 1.02 (0.93; 1.12) 0.90 (0.85; 0.96) (0; 1.06)
27 INSPIRE 1.00 0.97 (0.84; 1.12) 0.93 (0.90; 0.96) 0.93 (0.90; 0.96) (0; 1.01)
28 CAROLINA 1.30 0.98 (0.84; 1.14) 0.91 (0.79; 1.05) 0.94 (0.85; 1.05) (0; 1.05)
29 PRONOUNCE 1.00 1.28 (0.59; 2.79) 1.35 (0.94, 1.93) 1.34 (0.96; 1.85) (0; 1.85)
Table B.1: Confidence intervals (CIs) for the hazard ratio (HR) from each RCT and pooled RWE. Combined confidence intervals based on fixed effect meta-analysis and sceptical $p$-value are also reported, as well as the non-inferiority (NI) margin (1.00 for superiority studies).